OpenStudy (anonymous):

Can somebody check my answer? If ron can edit 15 essays in 10 hours and Dannielle needs 15 hours to do the same job. if they work together, how long will it take them to edit the essays?

OpenStudy (loser66):

where is your answer?

OpenStudy (loser66):

Ron: 15/10 = 3/2 =1.5 Dannielle 15/15 =1 that is , if they work together, Ron edits 3 essays, Dannielle can do 2, right? So the total essays is 15, Hence Ron does 9, and Daniell does 6 and they finish the work at the same time. NOW, Danielle does 15 essays / 15 hours , that is 1essay/ 1 hour he does 6essays and finishes at the same time with Ron. Hence, the total is........................

OpenStudy (anonymous):

@Loser66 the answer is 2. my answer is 2. the answer key says 6..

OpenStudy (anonymous):

is this right? im not sure how you did it, but i did it using rational equations (current unit we are in)

OpenStudy (loser66):

NO WAY to get 2.

OpenStudy (loser66):

the fastest guy is Ron, but he works in 10 hours to finish the stuff. If Danielle works as fast as Ron and they work together, they need 5 hours to finish the stuff. But Danielle is slower than Ron, he needs 15 hours to finish the same amount of Ron's work, Hence the answer CAN'T BE 2

OpenStudy (anonymous):

@Loser66 how do you know that ron can edit 3 hours?

OpenStudy (loser66):

when did I say Ron can edit 3 hours?